OpenStudy (anonymous):

Carlos and Maria drove a total of 233 miles in 4.4 hours. Carlos drove the first part of the trip and averaged 55 miles per hour. Maria drove the remainder of the trip and averaged 50 miles per hour. For approximately how many hours did Maria drive? Round your answer to the nearest tenth if necessary.

OpenStudy (anonymous):

Subscripts: c = carlos  and  м = maria      Dc + Dм = 233 mi ... distances      Dc = 233  –  Dм ... equation 1        Tc + Tм = 4.4 hrs ... times           Tc = 4.4  –  Tм ... equation 2  Distance = Velocity • Time     Dм = (Vм) • (Tм) ... Maria's equation     Dм = (50) • (Tм) ... equation 3             Tc = (Dc) ⁄ (Vc) ... equation 4 ... Carlos's equation             Tc = (Dc) ⁄ 55 ... substitute eqs. 1 & 2            4.4 – Tм = (233  –  Dм) ⁄ 55          55(4.4 – Tм) = 233  –  Dм ... substitute eq. 3          55(4.4 – Tм) = 233  –  50(Tм)         242  –  55Tм = 233  –  50(Tм)              5Tм = 9               Tм = 9 ⁄ 5 = 1.8 hours  ...  Maria's time